# S2AM3D: Scale-controllable Part Segmentation of 3D Point Cloud

### v1

@@ 5.3 Ablation Studies
Our Curated Data. To verify the impact of our own data, we train our model with PartNet [mo2019partnet] dataset, which results in performance drops compared to the model trained with our data. This finding strongly indicates that our curated data, having undergone systematic filtering and annotation governance, provides a crucial distributional complement to existing benchmarks. Specifically, it introduces richer geometric variation and diverse part composition, which directly translates into stable cross-dataset generalization gains for our model.
@@
Full 53.75 70.63 62.19
w/o 3D supervision 47.34 58.01 52.68
w/o our data 46.83 59.40 53.12
@@
Full 41.70 54.39 48.05
w/o 3D supervision 31.04 49.93 40.49
w/o our data 40.79 53.80 47.30
w/o scale embedding 40.31 53.28 46.80
@@ 6 Conclusion
We present S2AM3D, a scale-controllable framework for part segmentation of 3D point clouds. Through the synergistic design of a point-consistent encoder and a scale-aware prompt decoder, our method effectively addresses the challenge of cross-view inconsistency, enhances global feature consistency, and enables real-time adjustment of segmentation granularity via continuous scale signals. The construction of a large-scale, high-quality part-level point cloud dataset provides crucial support for model training. Extensive experiments demonstrate that S2AM3D achieves leading performance on multiple benchmarks, exhibiting exceptional robustness and controllability when handling complex structures and multi-scale parts, thereby providing a reliable solution for fine-grained 3D scene understanding and content editing. While the current framework primarily relies on point prompts and scale signals for interaction, future work will explore incorporating richer prompt modalities such as text instructions to support more intuitive semantic interaction.

### v3

@@ 5.3 Ablation Studies
Our Curated Data. To verify the impact of our own data, we train our model with PartNet [25] dataset, which results in performance drops compared to the model trained with our data. This finding strongly indicates that our curated data, having undergone systematic filtering and annotation governance, provides a crucial distributional complement to existing benchmarks. Specifically, it introduces richer geometric variation and diverse part composition, which directly translates into stable cross-dataset generalization gains for our model.
@@
Full 61.19 77.51 69.35
w/o 3D supervision 53.94 64.11 59.03
w/o our data 53.12 66.12 59.62
@@
Full 46.47 62.52 54.50
w/o 3D supervision 41.14 55.39 48.27
w/o our data 42.12 58.56 50.34
w/o scale embedding 42.31 58.28 50.30
@@ 6 Conclusion
We present S2AM3D, a scale-controllable framework for part segmentation of 3D point clouds. Through the design of a point-consistent encoder and a scale-aware prompt decoder, our method addresses the challenge of cross-view inconsistency, enhances global feature consistency, and enables real-time adjustment of segmentation granularity via continuous scale signals. The construction of a large-scale, high-quality part-level point cloud dataset provides crucial support for model training. Extensive experiments demonstrate that S2AM3D achieves leading performance on multiple benchmarks, exhibiting robustness and controllability when handling complex structures and multi-scale parts, thereby providing a reliable solution for fine-grained 3D scene understanding and content editing. While the current framework primarily relies on point prompts and scale signals for interaction, future work will explore incorporating richer prompt modalities such as text instructions to support more intuitive semantic interaction.
Acknowledgements
This work was supported by the National Key R&D Program of China under Grant No. 2022YFA1004100 and the National Natural Science Foundation of China (NSFC) under Grant No. 62476067.